Cupcake123456:

Jon invests $4,100 into an account that pays 5 percent simple interest each year. What is the total balance in the account after 3 years? $615.00 $3,485.00 $4,305.00 $4,715.00 @angle

Angle:

Here is an equation for simple interest that we can use (Deposit) * (number of years) * (1 + interest rate) = (Total Balance)

Angle:

so we are given the following information: Deposit = 4100 number of years = 3 interest rate = 0.05

Angle:

so first, what does the equation look like once we plug in all these numbers?

Cupcake123456:

do i mulitiply ?

Angle:

yes (4100)*(3)*(1.05) = ?

Cupcake123456:

12915?

Angle:

oh oh oh I got the equation wrong that was my fault there should not have been a (1) in the interest part

Angle:

so (4100)*(3)*(0.05) = interest then 4100 + interest = total

Cupcake123456:

4715.00 ?

Angle:

perfect :)

Cupcake123456:

*,